# Extracting memorized pieces of (copyrighted) books from open-weight language models

### v3

@@ E.5 Limits on “the monkey at the typewriter” metaphor
A common trope is to compare an LLM to a “monkey at the typewriter” [10, 125]. The basic idea is that, given infinite time, an LLM could presumably output any continuation of a piece of text. Like the monkey, it could by happenstance output a pristine copy of a Shakespeare play. This metaphor seems founded in the idea that LLMs are probabilistic. The monkey is randomly typing letters, and so, too, is an LLM, since there is an element of randomness in the generation process. For a given piece of text (the context), the LLM samples the next token to output based on the next-token conditional probabilities. That sampling is random, with respect to the conditional probability distribution of the possible next tokens.Already here, this description of what an LLM is doing to generate text should start to sound very different from a “monkey at the typewriter.” In that thought experiment, what the monkey is doing is random in the sense that every next character is equally likely to be produced at every step. But clearly, that is not what is happening with an LLM. The LLM has learned probabilistic relationships from the natural-language and other text (e.g., code) it was trained on. These relationships condition outputs given inputs. So while there is sampling involved—and thus probabilistic outcomes—those probabilities are not “random” in the same sense as the monkey typing. If an LLM could be used to produce a copy of a Shakespeare play (or Harry Potter and the Sorcerer’s Stone, Section 7 & Appendix F), it would not be because it behaves randomly like the monkey. (One could set the temperature to an enormous number, effectively inducing “monkey” behavior via the decoding scheme, but this is not how LLMs are used in practice.)
@@
  • If you stare at just the exact right part of the toothpick, and measure the length from the tip, expressed in terms of the appropriate unit and converted into binary, and then translated into English, you can find any message you want. You just have to pinpoint your gaze very very exactly (I call this “a prompt”). In fact, on your toothpick you can find the lead article from today’s New York Times. With enough squinting, measuring, and translating. By producing the toothpick, they put the message there and thus they gave you NYT access, even though you are not a paid subscriber. You simply need to know how to stare (and translate), or in other words how to prompt. So let’s sue the toothpick company! [33]
@@
There is plenty to say about misplaced metaphors about generative AI. We defer to others who have written extensively about this [76, 29, 28]. Here, we highlight how our specific results make specific points clear about why these two particular metaphors are both misguided. The key point is that our experiments show that (a) there will be at least some target suffixes that could never be generated by an LLM, even with infinite prompts to the model with a given prefix, and (b) a specific desired target suffix cannot always be produced by an LLM, given a sufficiently long prompt. The evidence for both comes from the same results: for a given LLM $\theta$ and decoding scheme $\phi$, some suffixes are impossible to generate, regardless of (a) how many times one independently prompts the model with a given prefix or (b) how long of a prefix one uses to condition the output.Our results make this very clear in practice, but we can also make this point conceptually. Intuitively, for a sequence $\bm{z}$}, if the model $\theta$ with decoding scheme $\phi$ assigns the verbatim suffix (given the prefix) zero probability (i.e., $p_{{\bm{z}}}{=}0$), then no amount of independent sampling trials will ever produce that suffix. If instead $p_{{\bm{z}}}>0$, then repeated independent sampling ensures that the probability of seeing $\bm{z}$} at least once tends to $1$ as the number of trials grows. Therefore, for any finite set of sequences $\mathbb{Z}$}, the maximum number of verbatim suffixes that can ever be generated is exactly those with nonzero probability:
@@
There are various ways that a given suffix could have $p_{\bm{z}}{=}0$. It might be the case that, for a particular decoding scheme, it is impossible to sample certain tokens, making some suffixes unreachable. This is the case for top-$k$ decoding, which only considers the top-$k$-highest probability tokens at each sampling iteration. Any token not included in the top-$k$ tokens forecloses possible output generation paths (Appendix A.3). For $k=40$ and for Llama 1 models, only $40$ out of 32,00032{,}000 tokens could be generated in a given sampling iteration. For another, numbers are represented in computers with a finite number of bits. This means that there is a minimum number that can be represented (>0>0) before it gets rounded down to 0. For instance, imagine a token vocabulary of 32,00032{,}000 tokens (as with Llama 1 models). If the model just generated tokens completing randomly (by setting temperature very high, like a “monkey at the typewriter”), then at each iteration every token would have probability $\frac{1}{32000}$. When generating $71$ such tokens in a row (i.e., a $71$-token suffix), the probability of any such suffix is ${\frac{1}{32000}}^{71}\approx 1.3626\times 10^{-320}$. Once we try to generate a 7272nd token, the probability for the whole sequence becomes ${\frac{1}{32000}}^{72}=0$. In terms of how the probability gets represented in the computer, there is underflow (i.e., rounding down to 0). As a result, even with completely-random temperature sampling, not every sequence could possibly be generated.We often observe 0-probability suffixes regardless of prefix length. This is true for non-training data (Section 5.2, Appendix E.4.2) and for training data that are not memorized (Section 5.1, Appendix E.3). Even if we add more context to statistically bias the output—to try to condition it in a particular direction—there are many cases where we simply are unable to raise $p_{\bm{z}}$ above 0 (let alone above $\tau_{\text{min}}$, the minimum probability we consider evidence for extraction).
@@ E.6.1 The relationship between membership inference and extraction
As noted in Carlini et al. [16], “Given a set of samples from the model, the problem of training data extraction reduces to one of membership inference: predict whether each sample was present in the training data.” This makes sense intuitively. Extraction is defined with respect to memorization: it is only possible to extract memorized training data, and (tautologically) training data can only be memorized if they are included—i.e., are members—of the training dataset. To demonstrate extraction is therefore to demonstrate memorization, and memorization implies membership. Not only does extraction imply membership of a piece of data, it actually reproduces that data in model outputs. In this respect, it is a stronger evidentiary claim (and, at least in principle, a harder one to make): membership inference asks a one-bit question—was this sequence a member of the training data?—while extraction produces the sequence itself. (At the same time, reliable membership inference with even the strongest membership inference attacks is extraordinarily challenging for LLMs [59], while extraction is relatively straightforward.) Put differently, one can make membership inference claims without extraction (e.g., with classifiers, see below); these claims can involve evidence where extraction might fail. So, even though successful extraction always implies membership, this does not go in both directions: membership does not imply successful extraction (Appendix E.6.5). Carlini et al. [16] goes on to note that “In their most basic form, past membership inference attacks rely on the observation that models tend to assign higher confidence to examples that are present in the training data. Therefore, a potentially high-precision membership inference classifier is to simply choose examples that are assigned the highest likelihood by the model.” At a high level, this is clearly directly related to extraction. Greedy-decoded discoverable extraction approximates the highest-likelihood sequence (albeit poorly) with greedy decoding [18]. Probabilistic extraction does the same but allows more flexibility than greedy decoding, and can identify sequences with higher likelihood than greedy-decoded ones [60].
@@ E.6.2 We know ground-truth membership for most of our experiments
In much work on extraction, researchers do not demonstrate extraction to infer membership. Instead, they run extraction experiments on known training data [60, 18, 74, 59], i.e., where ground-truth membership is not in doubt. This makes sense as a setup for drawing valid conclusions about extraction and memorization: membership is a prerequisite for memorization, and memorization is a prerequisite for extraction. Testing extraction on known member sequences therefore enables straightforward conclusions about memorization. Here, memorization is the most plausible explanation for extraction. As Carlini explains, when a sufficiently large and unique piece of training data is extracted, “the only possible explanation is that the model has somewhere internally stored [that piece of training data]” [15]. As Cooper and Grimmelmann similarly note, “in order to be able to extract memorized content from a model at generation time, that memorized content must be encoded in the model’s parameters. There is nowhere else it could be. A model is not a magical portal that pulls fresh information from some parallel universe into our own” [28, p. 25]. (See Section 2.)Most of our experiments fall into this category, as the large majority of our results are for Llama and Pythia models, where it is known that Books3 was included in the training data (Sections 46; Appendices D.1E.2). This is precisely the motivation behind our baseline experiments (Section 5.1, Appendix E.3), which we use as a basis of comparison for our negative controls (Section 5.2, Appendix E.4). Not only do we know that these models were trained on the books in Books3, we know that they were trained precisely on the Books3 corpus; we know the exact format of these data, just as they were presented to each of these models during training. Nevertheless, while work on greedy-decoded discoverable extraction does not do this, we perform validity experiments anyway. We need to set a minimum valid extraction probability $\tau_{\text{min}}$ (Appendix E.1). In this paper, we use a very high global $\tau_{\text{min}}$ to avoid false positives, and we validate this choice with our negative controls. Hayes et al. [60] do not set an explicit $\tau_{\text{min}}$, and do not make strong claims about when probabilistic extraction is valid; however, they do run a negative control experiment to show more generally that probabilistic extraction is a reasonable metric.In the original version of this paper, we included the negative controls on Phi 4 (Appendix E.4.1) to support our claims. Combined with the extensive analysis we perform on known training data for Llama, this is how we chose a global $\tau_{\text{min}}$: a conservative threshold that would not over-count extraction (and thus memorization). Even before including our experiments on post-cutoff books, our results provided ample evidence that non-training data exhibit 0 or orders of magnitude smaller probabilities $p_{\bm{z}}$ than training data. The post-cutoff negative controls provide further confirmation. This was true both for smaller, lower quality models, and higher quality models that exhibit lower loss (on average). While larger models like Llama 3.1 70B fit their training data better than smaller ones, and exhibit lower loss on average, they do not (and cannot) exhibit low loss everywhere, such that our measurements of $\log$ probabilities would be confounded. This is also clear from our experiments on known training data—Books3 books. For many books and large models, we are unable to extract any unique sequences; there are effectively no nonzero sequence probabilities. It is plausible that future work will find lower, model-specific $\tau_{\text{min}}$ thresholds, which may perhaps need to be set higher for larger models than for smaller ones. For now, however, the threshold we set is so conservative that these nuances do not affect our conclusions.
@@ E.6.3 We are inferring membership for models with undisclosed training data
In earlier versions of this paper, we placed our validity discussion in the appendix, but now highlight it in the main text (Section 5). Given our extensive results on known training data—including experiments to vet their validity—we can apply our methodology with confidence to models whose training data are not known with certainty. We do this for three model families: DeepSeek v1, Qwen 2.5, and Gemma 2. For these models, validated extraction serves as evidence of membership. These experiments differ from those on models known to be trained on Books3, but we ensure their validity through extensive negative controls and baselines. This approach is consistent with prior work. Carlini et al. [16] performed experiments on GPT-2. Nasr et al. [93, 94] extracted training data from open-weight models with undisclosed training data and from ChatGPT, a closed system. In both cases, the authors relied on proxy datasets and strategies to reduce bias in their results, such as shorter prompts. More generally, this is an important and reasonable thing to try to achieve. As researchers, we should be developing valid techniques for extraction that are useful in cases for which we do not know ground-truth membership. If membership inference required knowing membership in advance, there would be nothing to infer; at best—when it works—it would confirm what we already know. We do not need to use shorter prompts in our work because we validate our work with extensive negative controls. Further, we verified that every sequence extracted from Phi 4 and from post-cutoff books was widely duplicated online. Our validity experiments go beyond most prior work, which often does not include negative controls. This is intentional; our work has the potential to impact copyright litigation, not just the scientific literature. We therefore believe it is important to hold our work to a higher bar. Importantly, since training data for these models are undisclosed, we make claims only at the sequence level, not at the level of whole books (Appendix E.6.4). Finally, the population of data we study differs from most prior work, which evaluates memorization broadly across diverse training data (websites, code, spam, etc.) [16, 18, 93, 94, 60]. Such work must account for false positives like “trivial memorization” (e.g., number lists) or “repeated substrings” (e.g., “I love you” loops). LLMs assign high probability to both of these types of sequences, even if they are not memorized. We note these phenomena explicitly where we identify them, but they are generally not relevant our domain of long-form narrative text. The rare cases we identify of trivial memorization—lists of page numbers (an artifact of converting epub to txt), chapter indices—are excluded from our claims about extraction of unique creative expression. We identify one case of repeated substrings. Silverman [117] exhibits elevated probability in a part of the book that repeats “I will not wet the bed.” four times in a row. But this sequence does not surpass $\tau_{\text{min}}$, anyway, so we do not count it as extracted. (We also exclude boilerplate instances of true memorization from our claims, like copyright notices and author biographies.) Across more than one million sequences, these account for only a few hundred cases. This highlights why the population under study matters for validity: our focus on books yields different considerations than prior work, and we address these systematically. Unlike the work done in Carlini et al. [16], these types of edge cases are extremely rare in our population of interest.

### v4

@@ E.5 Limits on “the monkey at the typewriter” metaphor
A common trope is to compare an LLM to a “monkey at the typewriter” [borelmonkey, simpsons]. The basic idea is that, given infinite time, an LLM could presumably output any continuation of a piece of text. Like the monkey, it could by happenstance output a pristine copy of a Shakespeare play. This metaphor seems founded in the idea that LLMs are probabilistic. The monkey is randomly typing letters, and so, too, is an LLM, since there is an element of randomness in the generation process. For a given piece of text (the context), the LLM samples the next token to output based on the next-token conditional probabilities. That sampling is random, with respect to the conditional probability distribution of the possible next tokens.
Already here, this description of what an LLM is doing to generate text should start to sound very different from a “monkey at the typewriter.” In that thought experiment, what the monkey is doing is random in the sense that every next character is equally likely to be produced at every step. But clearly, that is not what is happening with an LLM. The LLM has learned probabilistic relationships from the natural-language and other text (e.g., code) it was trained on. These relationships condition outputs given inputs. So while there is sampling involved—and thus probabilistic outcomes—those probabilities are not “random” in the same sense as the monkey typing. If an LLM could be used to produce a copy of a Shakespeare play (or Harry Potter and the Sorcerer’s Stone, Section 7 & Appendix F), it would not be because it behaves randomly like the monkey. (One could set the temperature to an enormous number, effectively inducing “monkey” behavior via the decoding scheme, but this is not how LLMs are used in practice.)
@@
  • If you stare at just the exact right part of the toothpick, and measure the length from the tip, expressed in terms of the appropriate unit and converted into binary, and then translated into English, you can find any message you want. You just have to pinpoint your gaze very very exactly (I call this “a prompt”).
    In fact, on your toothpick you can find the lead article from today’s New York Times. With enough squinting, measuring, and translating.
    By producing the toothpick, they put the message there and thus they gave you NYT access, even though you are not a paid subscriber. You simply need to know how to stare (and translate), or in other words how to prompt.
    So let’s sue the toothpick company! [toothpick]
@@
There is plenty to say about misplaced metaphors about generative AI. We defer to others who have written extensively about this [lee2023talkin, cooper2023report, cooper2024files]. Here, we highlight how our specific results make specific points clear about why these two particular metaphors are both misguided. The key point is that our experiments show that (a) there will be at least some target suffixes that could never be generated by an LLM, even with infinite prompts to the model with a given prefix, and (b) a specific desired target suffix cannot always be produced by an LLM, given a sufficiently long prompt. The evidence for both comes from the same results: for a given LLM $\theta$ and decoding scheme $\phi$, some suffixes are impossible to generate, regardless of (a) how many times one independently prompts the model with a given prefix or (b) how long of a prefix one uses to condition the output.
Our results make this very clear in practice, but we can also make this point conceptually. Intuitively, for a sequence $\bm{z}$}, if the model $\theta$ with decoding scheme $\phi$ assigns the verbatim suffix (given the prefix) zero probability (i.e., $p_{{\bm{z}}}{=}0$), then no amount of independent sampling trials will ever produce that suffix. If instead $p_{{\bm{z}}}>0$, then repeated independent sampling ensures that the probability of seeing $\bm{z}$} at least once tends to $1$ as the number of trials grows. Therefore, for any finite set of sequences $\mathbb{Z}$}, the maximum number of verbatim suffixes that can ever be generated is exactly those with nonzero probability:
@@
There are various ways that a given suffix could have $p_{\bm{z}}{=}0$. It might be the case that, for a particular decoding scheme, it is impossible to sample certain tokens, making some suffixes unreachable. This is the case for top-$k$ decoding, which only considers the top-$k$-highest probability tokens at each sampling iteration. Any token not included in the top-$k$ tokens forecloses possible output generation paths (Appendix A.3). For $k=40$ and for Llama 1 models, only $40$ out of 32,00032{,}000 tokens could be generated in a given sampling iteration.
For another, numbers are represented in computers with a finite number of bits. This means that there is a minimum number that can be represented (>0>0) before it gets rounded down to 0. For instance, imagine a token vocabulary of 32,00032{,}000 tokens (as with Llama 1 models). If the model just generated tokens completing randomly (by setting temperature very high, like a “monkey at the typewriter”), then at each iteration every token would have probability $\frac{1}{32000}$. When generating $71$ such tokens in a row (i.e., a $71$-token suffix), the probability of any such suffix is ${\frac{1}{32000}}^{71}\approx 1.3626\times 10^{-320}$. Once we try to generate a 7272nd token, the probability for the whole sequence becomes ${\frac{1}{32000}}^{72}=0$. In terms of how the probability gets represented in the computer, there is underflow (i.e., rounding down to 0). As a result, even with completely-random temperature sampling, not every sequence could possibly be generated.
We often observe 0-probability suffixes regardless of prefix length. This is true for non-training data (Section 5.2, Appendix E.4.2) and for training data that are not memorized (Section 5.1, Appendix E.3). Even if we add more context to statistically bias the output—to try to condition it in a particular direction—there are many cases where we simply are unable to raise $p_{\bm{z}}$ above 0 (let alone above $\tau_{\text{min}}$, the minimum probability we consider evidence for extraction).
@@ E.6.1 The relationship between membership inference and extraction
As noted in carlini2021extracting, “Given a set of samples from the model, the problem of training data extraction reduces to one of membership inference: predict whether each sample was present in the training data.” This makes sense intuitively. Extraction is defined with respect to memorization: it is only possible to extract memorized training data, and (tautologically) training data can only be memorized if they are included—i.e., are members—of the training dataset. To demonstrate extraction is therefore to demonstrate memorization, and memorization implies membership.
Not only does extraction imply membership of a piece of data, it actually reproduces that data in model outputs. In this respect, it is a stronger evidentiary claim (and, at least in principle, a harder one to make): membership inference asks a one‐bit question—was this sequence a member of the training data?—while extraction produces the sequence itself. (At the same time, reliable membership inference with even the strongest membership inference attacks is extraordinarily challenging for LLMs [hayes2025strongmia], while extraction is relatively straightforward.) Put differently, one can make membership inference claims without extraction (e.g., with classifiers, see below); these claims can involve evidence where extraction might fail. So, even though successful extraction always implies membership, this does not go in both directions: membership does not imply successful extraction (Appendix E.6.5).
carlini2021extracting goes on to note that “In their most basic form, past membership inference attacks rely on the observation that models tend to assign higher confidence to examples that are present in the training data. Therefore, a potentially high-precision membership inference classifier is to simply choose examples that are assigned the highest likelihood by the model.” At a high level, this is clearly directly related to extraction. Greedy‐decoded discoverable extraction approximates the highest‐likelihood sequence (albeit poorly) with greedy decoding [carlini2023quantifying]. Probabilistic extraction does the same but allows more flexibility than greedy decoding, and can identify sequences with higher likelihood than greedy‐decoded ones [hayes2025measuringmemorizationlanguagemodels].
@@ E.6.2 We know ground-truth membership for most of our experiments
In much work on extraction, researchers do not demonstrate extraction to infer membership. Instead, they run extraction experiments on known training data [hayes2025measuringmemorizationlanguagemodels, carlini2023quantifying, lee2022dedup, hayes2025strongmia], i.e., where ground-truth membership is not in doubt. This makes sense as a setup for drawing valid conclusions about extraction and memorization: membership is a prerequisite for memorization, and memorization is a prerequisite for extraction. Testing extraction on known member sequences therefore enables straightforward conclusions about memorization. Here, memorization is the most plausible explanation for extraction. As Carlini explains, when a sufficiently large and unique piece of training data is extracted, “the only possible explanation is that the model has somewhere internally stored [that piece of training data]” [carlini2025blog]. As Cooper and Grimmelmann similarly note, “in order to be able to extract memorized content from a model at generation time, that memorized content must be encoded in the model’s parameters. There is nowhere else it could be. A model is not a magical portal that pulls fresh information from some parallel universe into our own” [cooper2024files, p. 25]. (See Section 2.)
Most of our experiments fall into this category, as the large majority of our results are for Llama and Pythia models, where it is known that Books3 was included in the training data (Sections 46; Appendices D.1 & E.2). This is precisely the motivation behind our baseline experiments (Section 5.1, Appendix E.3), which we use as a basis of comparison for our negative controls (Section 5.2, Appendix E.4). Not only do we know that these models were trained on the books in Books3, we know that they were trained precisely on the Books3 corpus; we know the exact format of these data, just as they were presented to each of these models during training.
Nevertheless, while work on greedy-decoded discoverable extraction does not do this, we perform validity experiments anyway. We need to set a minimum valid extraction probability $\tau_{\text{min}}$ (Appendix E.1). In this paper, we use a very high global $\tau_{\text{min}}$ to avoid false positives, and we validate this choice with our negative controls. hayes2025measuringmemorizationlanguagemodels do not set an explicit $\tau_{\text{min}}$, and do not make strong claims about when probabilistic extraction is valid; however, they do run a negative control experiment to show more generally that probabilistic extraction is a reasonable metric.
In the original version of this paper, we included the negative controls on Phi 4 (Appendix E.4.1) to support our claims. Combined with the extensive analysis we perform on known training data for Llama, this is how we chose a global $\tau_{\text{min}}$: a conservative threshold that would not over-count extraction (and thus memorization). Even before including our experiments on post-cutoff books, our results provided ample evidence that non-training data exhibit 0 or orders of magnitude smaller probabilities $p_{\bm{z}}$ than training data. The post-cutoff negative controls provide further confirmation.
This was true both for smaller, lower quality models, and higher quality models that exhibit lower loss (on average). While larger models like Llama 3.1 70B fit their training data better than smaller ones, and exhibit lower loss on average, they do not (and cannot) exhibit low loss everywhere, such that our measurements of $\log$ probabilities would be confounded. This is also clear from our experiments on known training data—Books3 books. For many books and large models, we are unable to extract any unique sequences; there are effectively no nonzero sequence probabilities. It is plausible that future work will find lower, model-specific $\tau_{\text{min}}$ thresholds, which may perhaps need to be set higher for larger models than for smaller ones. For now, however, the threshold we set is so conservative that these nuances do not affect our conclusions.
@@ E.6.3 We are inferring membership for models with undisclosed training data
In earlier versions of this paper, we placed our validity discussion in the appendix, but now highlight it in the main text (Section 5). Given our extensive results on known training data—including experiments to vet their validity—we can apply our methodology with confidence to models whose training data are not known with certainty. We do this for three model families: DeepSeek v1, Qwen 2.5, and Gemma 2. For these models, validated extraction serves as evidence of membership. These experiments differ from those on models known to be trained on Books3, but we ensure their validity through extensive negative controls and baselines.
This approach is consistent with prior work. carlini2021extracting performed experiments on GPT-2. nasr2023scalable, nasr2025scalable extracted training data from open-weight models with undisclosed training data and from ChatGPT, a closed system. In both cases, the authors relied on proxy datasets and strategies to reduce bias in their results, such as shorter prompts. More generally, this is an important and reasonable thing to try to achieve. As researchers, we should be developing valid techniques for extraction that are useful in cases for which we do not know ground-truth membership. If membership inference required knowing membership in advance, there would be nothing to infer; at best—when it works—it would confirm what we already know.
We do not need to use shorter prompts in our work because we validate our work with extensive negative controls. Further, we verified that every sequence extracted from Phi 4 and from post-cutoff books was widely duplicated online. Our validity experiments go beyond most prior work, which often does not include negative controls. This is intentional; our work has the potential to impact copyright litigation, not just the scientific literature. We therefore believe it is important to hold our work to a higher bar. Importantly, since training data for these models are undisclosed, we make claims only at the sequence level, not at the level of whole books (Appendix E.6.4).
@@
Finally, the population of data we study differs from most prior work, which evaluates memorization broadly across diverse training data (websites, code, spam, etc.) [carlini2021extracting, carlini2023quantifying, nasr2023scalable, nasr2025scalable, hayes2025measuringmemorizationlanguagemodels]. Such work must account for false positives like “trivial memorization” (e.g., number lists) or “repeated substrings” (e.g., “I love you” loops). LLMs assign high probability to both of these types of sequences, even if they are not memorized. We note these phenomena explicitly where we identify them, but they are generally not relevant our domain of long-form narrative text. The rare cases we identify of trivial memorization—lists of page numbers (an artifact of converting epub to txt), chapter indices—are excluded from our claims about extraction of unique creative expression. We identify one case of repeated substrings. bedwetter exhibits elevated probability in a part of the book that repeats “I will not wet the bed.” four times in a row. But this sequence does not surpass $\tau_{\text{min}}$, anyway, so we do not count it as extracted. (We also exclude boilerplate instances of true memorization from our claims, like copyright notices and author biographies.) Across more than one million sequences, these account for only a few hundred cases. This highlights why the population under study matters for validity: our focus on books yields different considerations than prior work, and we address these systematically. Unlike the work done in carlini2021extracting, these types of edge cases are extremely rare in our population of interest.